x=5 + 3y how do you solve this?
10=4x + 2y how do you solve this?
y=4-3x ??

Answer by rfer(16322)

solve for ordered pairs, make a table and graph.
y=f(x)=(x-5)/3
y=f(x)=(-4x+10)/2
y=f(x)=-3x+4